# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation

### v3

@@ 6 Empirical Experiments
We conducted NLG tests on Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023). We trained the models with SORSA applied using MetaMathQA (Yu et al., 2024), and evaluated the model on GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). We also trained the model on CodeFeedback Filtered Instruction (Zheng et al., 2024) dataset and evaluated it on HumanEval (Chen et al., 2021). The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024). All reported values are accuracy. See Appendix A.2 for more details and hyperparameters of the training. We quoted results for PiSSA, LoRA, and full parameter fine-tuning directly from Meng et al. (2024). For SORSA experiments, we conducted our experiments on an NVIDIA A100-SXM4 (80GB) GPU. See Table 1 for the results and Figure 3 for the loss and gradient norm comparison.
Model Method Trainable Parameters MATH GSM-8K HumanEval
Llama 2 7B Full FT 6738M 7.22 49.05 21.34
LoRA 320M 5.50 42.30 18.29
PiSSA 320M 7.44 53.07 21.95
SORSA 320M 10.36 56.03 24.39
Mistral 7B Full FT 7242M 18.60 67.02 45.12
LoRA 168M 19.68 67.70 43.90
PiSSA 168M 21.54 72.86 46.95
SORSA 168M 21.86 73.09 47.56
@@
Figure 3: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of Llama 2 7B. LoRA and PiSSA curves are from Meng et al. (2024).
The results showed that on the Llama 2 7B model, SORSA achieved 10.36% accuracy on the MATH benchmark and 56.03% accuracy on the GSM-8K benchmark, both significantly outperforming PiSSA, LoRA, and Full FT. On the Mistral 7B v0.1 model, SORSA achieved 21.86% accuracy on the MATH benchmark and 78.03% accuracy on the GSM-8K benchmark, demonstrating slightly better performance on both evaluation tasks compared to the other three methods.
In this section, due to the limitation of computing resources, we train and benchmark a small number of tasks.
@@ 7 Conclusion
We adopt singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis.
Our gradient analysis provides a mathematical foundation for SORSA, demonstrating its convexity, Lipschitz continuity, and the crucial role of the regularizer in improving the optimization landscape. This theoretical framework not only explains SORSA’s empirical superior performance but also offers valuable insights for future developments in adaptive learning algorithms.
@@ 8 Future Work
  • Extended Evaluation: Conduct a broader evaluation of SORSA across various benchmarks in different models, in order to provide a more comprehensive understanding of SORSA’s performance and versatility.
@@
  • Quantization: Investigate the application of quantization techniques to SORSA similar to the approach of QLoRA (Dettmers et al., 2024). This research aims to reduce the memory footprint and computational requirements of SORSA, making it feasible to deploy on edge devices or in environments with limited computational resources. By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, QSORSA could enable highly efficient and powerful language models suitable for a broader range of practical applications.
@@ 9 Impact Statement
In this paper, we introduce an innovative PEFT method in the area of Machine Learning. Our approach significantly streamlines the model’s tuning process, particularly for large-scale models, addressing both computational efficiency and environmental sustainability. As we push the boundaries of what is possible with machine learning, it is essential to consider the broader impacts of these advancements on both the environment and ethical standards within the field.
@@ 9.1 Environmental Impact
SORSA optimizes model tuning for large-scale models, reducing VRAM consumption by over 50% when adapting the Llama 3 70B model than full parameter fine-tuning. This significant reduction in hardware resource requirements translates to approximately 50% less energy consumption compared to full parameter fine-tuning methods. By enhancing efficiency, our approach contributes to reducing the environmental footprint of machine learning operations.
@@ 10 Acknowledgment
Appendix A Experiments Details
A.1 Analysis
For the singular values and vectors analysis in Section 3, we applied partial fine-tuning, LoRA and SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset. In the analysis, LoRA and SORSA were only applied to q_proj and v_proj matrices, respectively. For Partial FT, we set model’s q_proj and v_proj matrices to trainable. We only calculated the loss on the response part. The models are trained with TF32 + BF16 mix precision. See Table 2 for hyperparameters.
@@ A.1 Analysis
Method FT LoRA SORSA (w/o reg) SORSA
Precision TF32 + BF16 Mix-Precision
Epoch 1
Batch Size 128
Max Length 512
Weight Decay 0.00
Warm-up Ratio 0.03
@@
$\gamma$ 0 5e-4
Rank 128 128 128
@@
A.2 NLG Experiments
For our NLG tasks, we adapted Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by SORSA. For GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) evaluations, we trained those models with the first 100K data in MetaMathQA (Yu et al., 2024) dataset. For HumanEval (Chen et al., 2021) evaluation, we use the first 100K data in CodeFeedback Filtered Instruction(Zheng et al., 2024) dataset.
We used AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA adapters were applied on all linear matrices in every layer. We only calculated the loss on the response part. The models are trained with TF32 + BF16 mix precision. See Table 3 for hyperparameters.
@@ A.2 NLG Experiments
Model Llama 2 7B Mistral 7B v0.1
Method SORSA
Precision TF32 + BF16 Mix-Precision
Epoch 1
Batch Size 128
Max Length 512
Weight Decay 0.00
Warm-up Ratio 0.03
Learning Rate 3e-5
$\gamma$ 4e-4
Rank 128 64
Inference
Precision BF16
@@
GSM-8K Max Length 1024
MATH Max Length 2048
HumanEval Max Length 2048
@@
Listing 1: Prompt used for HumanEval evaluation.
Appendix B Proofs
B.1 Proof of Theorem 5.1
@@ Appendix B Proofs
We can assume that the regularizer’s gradients scale with $\eta_{d}$ , meaning that larger updating step (due to a larger $\eta_{d}$) will lead to greater deviations from orthonormality, which increases $L_{reg}$. Conversely, smaller steps lead to a more gradual progression towards orthonormality, which reduces $L_{reg}$. Therefore, we could assume $L_{reg}\propto\eta_{d}$, and rewrite the inequality as:
@@
$\nabla_{W_{p}}{\mathcal{L}}_{reg}\leq\epsilon_{\nabla}$ (23)
@@
$\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}=\gamma\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}$ (24)
Now, we can use Weyl’s inequality, which states that for matrices A and B:
@@
Let $W_{p}=U_{p}\Sigma_{p}V_{p}^{\top}$ be the principal part of the singular value decomposition approximation of $W$. The condition number is given by:
@@
$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=\frac{\kappa(\Sigma_{p}^{\text{reg}})}{\kappa(\Sigma_{p}^{\text{unreg}})}=1$ (31)
@@
From the Lemma 5.4, we arrive at:

### v4

@@ 6 Empirical Experiments
We conducted NLG tests on Llama 2 7B (Touvron et al., 2023), RWKV6 7B (Peng et al., 2024), Mistral 7B v0.1 (Jiang et al., 2023) Gemma 7B (Gemma Team et al., 2024). We trained the models using first 100K data in MetaMathQA (Yu et al., 2024), and evaluated the model on GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). We also trained the model on first 100K data in CodeFeedback Filtered Instruction (Zheng et al., 2024) dataset and evaluated it on HumanEval (Chen et al., 2021). The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024). All reported values are accuracy. See  Section B.2 for more details and hyperparameters of the training. We quoted some results of PiSSA, LoRA, and full parameter fine-tuning from Meng et al. (2024). For experiments, we conducted part of our experiments on a single NVIDIA A100-SXM4 (80GB) GPU, others on a single NVIDIA H100-SXM4 (80GB) GPU. See Table 1 for the results and Figure 3 for the loss and gradient norm comparison.
Model Method Trainable Parameters GSM-8K MATH HumanEval
Llama 2 7B Full FT 6738M 49.05 7.22 21.34
LoRA 320M 42.30 5.50 18.29
PiSSA 320M 53.07 7.44 21.95
SORSA 320M 56.03 10.36 24.39
RWKV6 7B LoRA 176M 8.04 7.38 15.24
PiSSA 176M 32.07 9.42 17.07
SORSA 176M 45.87 11.32 22.56
Mistral 7B Full FT 7242M 67.02 18.60 45.12
LoRA 168M 67.70 19.68 43.90
PiSSA 168M 72.86 21.54 46.95
SORSA 168M 73.09 21.86 47.56
Gemma 7B Full FT 8538M 71.34 22.74 46.95
LoRA 200M 74.90 31.28 53.66
PiSSA 200M 77.94 31.94 54.27
SORSA 200M 78.09 29.52 55.49
@@
Refer to caption Refer to caption
Figure 3: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of RWKV6 7B and Llama 2 7B. LoRA and PiSSA curves of Llama 2 7B are from Meng et al. (2024).
The results showed that across all models tested, SORSA generally outperformed other methods. For mathematical evaluations, on Llama 2 7B, SORSA scored 56.03% on GSM-8K and 10.36% on MATH, significantly outperforming other methods; For the RWKV6 7B model, SORSA achieved 45.87% accuracy on GSM-8K and 11.32% on MATH, surpassing both PiSSA and LoRA; On Mistral 7B, SORSA reached 73.09% on GSM-8K and 21.86% on MATH, slightly edging out competitors; With Gemma 7B, while SORSA’s MATH score of 29.52% was lower than LoRA and PiSSA, it still led on GSM-8K with 78.09%. Moreover, on coding evaluations, SORSA consistently achieved the highest HumanEval scores across other tuning methods, demonstrating its effectiveness on diverse natural language tasks. Noticeably, RWKV6 7B using LoRA performed extremely below our expectation in GSM-8K test. We believe this may be due to our prompt for mathematical tests (See Section B.2 for more details) are inconsistent with the model’s original prompt, causing LoRA to fail in learning this prompting behavior.
The Figure 3 reveals that while SORSA and PiSSA exhibit nearly identical loss curves at the beginning, and even a little bit higher than PiSSA on RWKV-6 training. However, when training step approximately $t>300$, SORSA continues to maintain the decrease of its loss steadily. In contrast, LoRA and PiSSA shows a deceleration in its loss reduction. This supports Theorem 5, confirming that SORSA consistently outperforms both LoRA and PiSSA, especially at later stages of training.
However, due to the limitation of computing resources, we only trained and benchmarked a small number of tasks.
@@ 7 Conclusion
We adopted singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis.
Our gradient analysis provided a mathematical foundation for SORSA, demonstrating its convexity, Lipschitz continuity, and the crucial role of the regularizer in improving the optimization landscape. This theoretical framework not only explains SORSA’s empirical superior performance but also offers valuable insights for future developments in adaptive learning algorithms.
@@ 8 Future Work
  • Comparison Between Different Rank: Find how the performance of SORSA varies with different ranks. This investigation would help determine the optimal rank for various tasks and model sizes, potentially revealing a trade-off between computational efficiency and model performance.
  • Extended Evaluation: Conduct broader evaluations of SORSA across different areas of benchmarks on models with more diverse sizes. This may able to provide a more comprehensive understanding of SORSA’s performance and versatility.
@@
  • Quantization: Investigate the application of quantization techniques to SORSA similar to the approach of QLoRA (Dettmers et al., 2024) and QPiSSA (Meng et al., 2024). Quantization aims to reduce the memory footprint and computational requirements, making it feasible to deploy on devices with limited computational resources. By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, quantization of SORSA may able to efficiently provides powerful models, which could eventually facilitate AI for a broader range of practical applications.
@@ 9 Impact Statement
In this paper, we introduced an innovative PEFT method in the area of Machine Learning. Our approach significantly streamlined the model’s tuning process, particularly for large-scale models, addressing both computational efficiency and environmental sustainability. As we push the boundaries of what is possible with Machine Learning, it is essential to consider the broader impacts of these advancements on both the environment and ethical standards within the field.
@@ 9.1 Environmental Impact
SORSA optimizes model tuning for large-scale models, reducing VRAM consumption by over 50% when adapting the Llama 3 70B Dubey et al. (2024) model than full parameter fine-tuning. This significant reduction in hardware resource requirements also suggests a less energy consumption compared to full parameter fine-tuning methods. By enhancing efficiency, our approach could significantly contributes to reducing the carbon footprint of Machine Learning operations.
@@ 10 Acknowledgment
Appendix A Faster SORSA Adapters
According to the definition of SORSA from Equation 7, because $\text{diag}(S_{p})$ is always a diagonal matrix, it is equivalent to:
$\text{SORSA}(x)=xW_{r}+x(U_{p}\odot S_{p})V_{p}^{\top}$ (14)
Where $\odot$ denotes element-wise multiplication, and $\text{diag}(W)$ denotes the diagonal elements of the matrix $W$.
This transformation allows us to reduce the computational complexity of SORSA adapters. In the original form, we have to perform twice matrix multiplications. However, in the Equation 14, we only have one matrix multiplication, and one element-wise multiplication. The time complexity of element-wise multiplication is $O(m^{2})$, which is much less than matrix multiplication’s complexity $O(m^{3})$. Therefore, while $\lim_{m\to\infty}$, the computation speed of SORSA adapters will be same as LoRA and PiSSA.
We performed a benchmark using PyTorch (Paszke et al., 2019) on a NVIDIA H100 SXM4 (80GB) GPU backed with CUDA, and Apple M2Pro CPU, in order to test the computation time between these two methods. See Figure 4 to see our results.
Refer to caption Refer to caption
Figure 4: Benchmark between two equations of SORSA
Appendix B Experiments Details
B.1 Analysis
For the singular values and vectors analysis in Section 3, we applied partial fine-tuning, LoRA and SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset. We only calculated the loss on the response part. The models are trained with TF32 & BF16 (Wang & Kanwar, 2019) mix precision. See Table 2 for hyperparameters.
We used AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. In the analysis, LoRA and SORSA were only applied to q_proj and v_proj matrices, respectively. For Partial FT, we set model’s q_proj and v_proj matrices to trainable.
@@ B.1 Analysis
Model Llama 2 7B
Method Partial FT LoRA SORSA (w/o reg) SORSA
@@
Mix-Precision TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16
Epoch 1 1 1 1
Batch Size 128 128 128 128
Max Length 512 512 512 512
Weight Decay 0 0 0 0
Warm-up Ratio 0.03 0.03 0.03 0.03
@@
Grad Clip False False False False
SORSA $\gamma$ N/A N/A 0 5e-4
Rank N/A 128 128 128
@@
B.2 NLG Experiments
For our NLG tasks, we adapted Llama 2 7B (Touvron et al., 2023), RWKV6 7B (Peng et al., 2024), Mistral 7B v0.1 (Jiang et al., 2023) Gemma 7B (Gemma Team et al., 2024). models by SORSA. For GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) evaluations, we trained those models with the first 100K data in MetaMathQA (Yu et al., 2024) dataset. For HumanEval (Chen et al., 2021) evaluation, we use the first 100K data in CodeFeedback Filtered Instruction(Zheng et al., 2024) dataset.
We used AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA adapters were applied on all linear matrices in every layer. We only calculated the loss on the response part. The models are loaded in FP32, and trained with TF32 & BF16 mix precision. In our experiments, we selected a higher learning rate for SORSA compared to other methods to counterbalance the negative effect of orthonormal regularizer on optimizing toward lower training loss. See Table 3 for hyperparameters. See LABEL:lst:math for the prompt we used in GSM-8K and MATH evaluations, and LABEL:lst:humaneval for the prompt we used for HumanEval tests.
Model Llama 2 7B Mistral 7B Gemma 7B RWKV6 7B RWKV6 7B
Method SORSA SORSA SORSA SORSA LoRA PiSSA
@@ B.2 NLG Experiments
Mix-Precision TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16
Epoch 1 1 1 1 1
Batch Size 128 128 128 128 128
Max Length 512 512 512 512 512
Weight Decay 0 0 0 0 0
Warm-up Ratio 0.03 0.03 0.03 0.03 0.03
Learning Rate 3e-5 3e-5 3e-5 3e-5 2e-5
Grad Clip 1.0 1.0 1.0 1.0 1.0
SORSA $\gamma$ 4e-4 4e-4 4e-4 4e-4 N/A
Rank 128 64 64 64 64
Evaluating
Precision BF16 BF16 BF16 FP32 FP32
Sampling False
@@
Max Length GSM-8K: 1024 MATH: 2048 HumanEval: 2048
@@
1Below is an instruction that describes a task. Write a response that appropriately completes the request.
2
3### Instruction:
4{question}
5
6### Response: Lets think step by step.
Listing 1: Prompt used for GSM-8K and MATH.
@@
Listing 2: Prompt used for HumanEval evaluation.
Appendix C Proofs
See 1
@@ Appendix C Proofs
We can assume that the regularizer’s gradients scale with $\eta_{d}$ , meaning that larger updating step (due to a larger $\eta_{d}$) will lead to greater deviations from orthonormality, which increases $L_{reg}$. Conversely, smaller steps lead to a more gradual progression towards orthonormality, which reduces $L_{reg}$. Therefore, we could assume $L_{reg}\propto\eta_{d}$. Moreover, the $\gamma$ must not be negative or the regularization term would impact negatively on the purposes it is supposed to have. Therefore, we can rewrite the inequality as:
@@
$\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}\leq\epsilon_{\nabla}$ (24)
@@
$\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}=\gamma\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}=\gamma\epsilon_{\nabla}$ (25)
Now, we can use Weyl’s inequality (Weyl, 1912), which states that for matrices A and B:
@@
Let $W_{p}=U_{p}\text{diag}(S_{p})V_{p}^{\top}$ be the principal part of the singular value decomposition approximation of $W$. The condition number is given by:
@@
$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=\frac{\kappa(\text{diag}(S_{p})^{\text{reg}})}{\kappa(\text{diag}(S_{p})^{\text{unreg}})}=1$ (32)
@@
From the Lemma 4, we arrive at: